OpenStudy (anonymous):

what is the integral od sin2x/csc^2 X sec^2 X ? ..ty :D

OpenStudy (anonymous):

I'm assuming you're saying \[\int\frac{\sin 2x}{\csc^2x\sec^2x}~dx\] Use the double angle identity to rewrite the numerator: \[\int\frac{2\sin x\cos x}{\csc^2x\sec^2x}~dx\] Since sine/cosecant and cosine/secant are reciprocal function, you have \[2\int\sin^3 x\cos^3 x~dx\] Reduce the power of either sine or cosine. I'll do it with the cosine: \[2\int\sin^3 x\cos^2 x\cos x~dx\] Pythagorean identity: \[2\int\sin^3 x(1-\sin^2x)\cos x~dx\] Distribute \[2\left(\int\sin^3 x\cos x~dx-\int\sin^2x\cos x~dx\right)\] Solve each integral with the substitution \(u=\sin x\).